a manufacturer knows that if x goods are demanded on a particular week, the total cost and revenue function will be: c[x]=14+3x and R[x]=18x-2x^2 respectively. calculate the level of demand that will maximize profits. calculate the amount of profit that will be realized at this maximum point


1
Expert's answer
2021-11-25T15:51:00-0500

i)

profit function:


"P(x)=\ud835\udc45(\ud835\udc65)-\ud835\udc36(\ud835\udc65)=18x-2x^2-(14+3x)"


"=-2x^2+15x-14, x\geq 0"

Differentiare with respect to "x"


"P'(x)=(-2x^2+15x-14)=-4x+15"

Find the critical number(s)


"P'(x)=0=>-4x+15=0=>x=15\/4"

If "0\leq x\leq15/4, P'(x)>0, P(x)" increases.


If "x>15\/4, P'(x)<0, P(x)" decreases.

The profit function "P(x)" has a local maximum at "x=15\/4."

Sinse the profit function "P(x)" has the only extremum, then the function "P(x)" has the absolute maximum at "x=15\/4."

Level of demand that will maximize profits:


"x=15\/4=3.75"


If "x" is an integer, then


"P(3)=-2(3)^2+15(3)-14=13"

"P(4)=-2(4)^2+15(4)-14=14"

We take "x=4."


ii)

The amount of profit that will be realized at this maximum point:


"P_{max}(x)=P(3.75)"




"=-2\cdot(3.75)^2+15\cdot3.75-14=14.125"

If "x" is an integer, then


"P_{max}(x)=P(4)"




"=-2\cdot(4)^2+15\cdot4-14=14"